Question: Find the slope and y-intercept of the line given by 2x+3y=5. I don't know how to work this. Is it possible to do on a TI-83 calculator? If so, how?
Answer: Slope=_________ y-intercept=____________

Found 2 solutions by Alan3354, fcabanski:
Answer by Alan3354(69443)

Find the slope and y-intercept of the line given by 2x+3y=5.
--------
Solve for y. That's all there is to it.
2x+3y=5
y = (-2/3)x + 5/3
The coeff of x is the slope, m
m = -2/3
--------
The y-int is the coeff, b
b = 5/3
=========
Solve for y, then it's the form y = mx + b
-----
I have no idea about that calculator, but it doesn't seem to be worth the effort for this one.
Someone is sending me a TI-83, but I haven't gotten it yet.

Answer by fcabanski(1391)

You can graph it on the calculator, but that won't make the slope apparent. Look it up with a google search.


The easiest way to find the slope and y-intercept of a line is to place it in the slope intercept form y=mx+b where m is the slope and b is the y intercept.


2x+3y=5 | Subtract 2x from both sides.
3y = -2x + 5 | divide everything by 3.
y = (-2/3)x + 5/3


The slope is -2/3 and the y intercept is 5.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com